# Is magnitude ‘generically continuous’ for finite metric spaces?

### v1

@@ 1. Introduction
Under favourable conditions—for instance, if $X$ is a subset of Euclidean space—this function is in fact continuous and monotonically increasing with $t$, and satisfies $\lim_{t\to 0}|tX|=1$ and $\lim_{t\to\infty}|tX|=\#X$. In such cases, the magnitude function can be interpreted as recording the ‘effective number of points’ in $X$ as the scale of the metric varies, and its instantaneous rate of growth as recording the ‘effective dimension’. For illustrative examples, see [19, Ex. 6.4.6] and [29, §4].
The definition of magnitude can be extended from finite metric spaces to a large class of compact metric spaces [22, §2], and in that setting it turns out to be remarkably rich in geometric content. The large-scale asymptotics of the magnitude function are of particular interest when interpreted for domains in Euclidean space [5, 9, 30, 23] or manifolds with boundary [11, 10], capturing information about ‘size-related’ attributes such as volume, surface area, and integrals of curvatures. There is also an associated homology theory, magnitude homology, which has been related in valuable ways to graph-theoretic invariants arising in the context of discrete homotopy theory [14, 4, 13, 16].
@@
is continuous (or, better, Lipschitz continuous). Unfortunately, magnitude is not continuous in this sense: examples of discontinuities can be found in [18, Ex. 2.2.7 and Ex. 2.2.8] and [27, Ex. 3.4]. Yet, such examples seem to be rare—so rare, in fact, that intuition suggests one is vanishingly unlikely to encounter instances of magnitude’s instability in ‘the wild’ of real-world data.
In this paper we take the point of view that, for the purposes of applicability, the generic behaviour of invariants is what one needs to understand, not their vanishingly rare pathologies. This seems to reflect actual practice in the domains where magnitude is being applied: for instance, in [21, §3.2], the assumption that $\lim_{t\to 0}|tX|=1$ is justified by the fact, proved in [27, Thm. 2.3], that this property holds generically—though not always—for finite metric spaces.
@@
Section 2 also contains new examples illustrating just how ‘badly’ continuity can fail. In [27, Thm. 3.8] it was shown that for every real number $\alpha\geq 1$, there exists a finite metric space $X$ such that $\lim_{t\to 0}|tX|=\alpha$, but the question was left open of whether there exists $X$ such that $\lim_{t\to 0}|tX|<1$ or $\lim_{t\to 0}|tX|=+\infty$. That question is not answered here, but in Example 2.6 we show that if one considers arbitrary paths $(X_{s})_{0<s\ll 1}$ in $\operatorname{FMet}^{\circ}$ converging to the one-point space, then $\lim_{s\to 0}|X_{s}|$ can take any value in the extended real line. We also consider certain paths converging to a two-point space $K_{2}$, and describe the associated sequences of magnitude functions. We demonstrate that the pointwise limit of these functions need not display any of the good qualitative features of the magnitude function of $K_{2}$, such as continuity or monotonicity with respect to the scale factor $t$, or meaningful behaviour in the large- or small-scale limit.
@@
An alternative route to a stability theorem for magnitude would be to look for a class of metric spaces with the property that if $X$ belongs to this class, then magnitude is continuous with respect to the Hausdorff topology on finite subsets of $X$. One good candidate might be the class of positive definite spaces. In [22], Meckes defines a positive definite metric space to be one such that every finite subspace has positive-definite similarity matrix, and he proves that magnitude is lower semicontinuous—though not continuous—on the Gromov–Hausdorff space of compact positive definite metric spaces [22, Thm. 2.6]. He also poses the following question:
@@
Since the metric induced by the $\ell^{p}$-norm on $\mathbb{R}^{n}$ is positive definite for $1\leq p\leq 2$ [22, Thm. 3.6], a positive answer to 1.2 would imply in particular that magnitude is continuous for finite subsets of these spaces. This would be especially valuable to know for the purposes of applications, which often concern data embedded into $\mathbb{R}^{n}$ with the Euclidean ($\ell^{2}$) or taxicab ($\ell^{1}$) metric. Moreover, in the setting of $\mathbb{R}^{n}$ the techniques of Leinster and Meckes [20] are available, and these seem likely to facilitate stronger results on continuity.
We therefore close this introduction by stating a conjecture.
@@
If $U=(\mathbb{R}^{n},\ell^{p})$ where $1\leq p\leq 2$, then the function $X\mapsto|X|$ is continuous with respect to the Hausdorff topology on the set of finite subsets of $U$. In particular, magnitude is stable for finite subsets of Euclidean or taxicab space.
@@ Acknowledgements
This work was partially supported by JSPS KAKENHI JP22K18668, JP23H00081. We are grateful to Sergei O. Ivanov for useful conversations.
@@ 2. Magnitude is nowhere continuous on Gromov–Hausdorff space
    [27, Example 3.4] Let $X=\frac{4}{3}K_{2}+_{1}2K_{3}$. Then $\lim_{t\to 0}|tX|=\nicefrac{{7}}{{6}}$.
@@
Take any $X\in\operatorname{FMet}^{\circ}$. Let $Y$ be a space without the one-point property—for concreteness, we may choose $Y=K_{3}+_{1}2K_{3}$ as in Example 2.4 (1). Pick any points $x\in X$ and $y\in Y$, take $t>0$ and consider the wedge $X\vee tY=X\vee_{x,y}tY$. Observe that $X$ embeds isometrically into $X\vee tY$, and that in the Hausdorff metric on subspaces of the wedge, we have $d(X,X\vee tY)\leq\operatorname{diam}(tY)\to 0$ as $t\to 0$. It follows that $X\vee tY\to X$ in $\operatorname{FMet}$. Meanwhile, for sufficiently small $t>0$ the space $tY$ has magnitude [27, Rmk. 2.1], so, by Lemma 2.3 (1), for sufficiently small $t>0$ the space $X\vee tY$ has magnitude given by $|X\vee tY|=|X|+|tY|-1$. Hence, as $t\to 0$, we have
@@
In [27, Thm. 3.8], it was shown that magnitude’s failure of continuity at the one-point space can be ‘arbitrarily bad’, in the sense that for each real number $\alpha\geq 1$ there exists a finite metric space $X$ with $\lim_{t\to 0}|tX|=\alpha$. So far as we know, no example has yet been found of a finite metric space $X$ such that $\lim_{t\to 0}|tX|<1$ or $\lim_{t\to 0}|tX|=+\infty$. However, if we consider more general paths in $\mathrm{FMet}$ converging to the one-point space, the limiting magnitude can attain any value in the extended reals. A modification of Example 2.4 realizes such paths.
@@
By [27, Thm. 3.8], for every real number $\alpha\geq 1$ there exists a finite metric space $Y_{\alpha}$ such that $\lim_{t\to 0}|tY_{\alpha}|=\alpha$. Fix $\alpha$ and let $W_{t}=K_{2}\vee tY_{\alpha}$; then $\lim_{t\to 0}W_{t}=K_{2}$. For each $s>0$, we have $sW_{t}=sK_{2}\vee stY_{\alpha}$. It follows, by Lemma 2.3 (1), that
@@ 3. Lines in Gromov–Hausdorff space
In [27, §2], the small-scale limit of the magnitude function was studied via this identification of finite metric spaces with equivalence classes of matrices. To adapt the techniques of that paper to the present setting, we will identify lines in Gromov–Hausdorff space with certain equivalence classes of block matrices.
@@ 4. Magnitude behaves continuously for generic limits along lines
hence, $\det(Z_{f,t})$ is a generalized polynomial in $e^{-t}$, in the sense of [27, §2] (but now with real coefficients). It follows that $\det(Z_{f,t})$ vanishes for at most finitely many $t$. In particular, for all sufficiently small $t>0$ we must have $\det(Z_{f,t})\neq 0$, so the expression (4.2) is valid for $0<t\ll 1$. We will compute the denominator and the numerator of (4.2) separately, beginning with the denominator.

### v2

@@ 1. Introduction
Under favourable conditions—for instance, if $X$ is a subset of Euclidean space—this function is in fact continuous, and satisfies $\lim_{t\to 0}|tX|=1$ and $\lim_{t\to\infty}|tX|=\#X$. In such cases, the magnitude function can be interpreted as recording the ‘effective number of points’ in $X$ as the scale of the metric varies, and its instantaneous rate of growth as recording the ‘effective dimension’. For illustrative examples, see [19, Ex. 6.4.6] and [31, §4].
The definition of magnitude can be extended from finite metric spaces to a large class of compact metric spaces [23, §2], and in that setting it turns out to be remarkably rich in geometric content. The large-scale asymptotics of the magnitude function are of particular interest when interpreted for domains in Euclidean space [5, 9, 32, 25] or manifolds with boundary [11, 10], capturing information about ‘size-related’ attributes such as volume, surface area, and integrals of curvatures. There is also an associated homology theory, magnitude homology, which has been related in valuable ways to graph-theoretic invariants arising in the context of discrete homotopy theory [14, 4, 13, 16].
@@
is continuous (or, better, Lipschitz continuous). Unfortunately, magnitude is not continuous in this sense: examples of discontinuities can be found in [18, Ex. 2.2.7 and Ex. 2.2.8] and [29, Ex. 3.4]. Yet, such examples seem to be rare—so rare, in fact, that intuition suggests one is vanishingly unlikely to encounter instances of magnitude’s instability in ‘the wild’ of real-world data.
In this paper we take the point of view that, for the purposes of applicability, the generic behaviour of invariants is what one needs to understand, not their vanishingly rare pathologies. This seems to reflect actual practice in the domains where magnitude is being applied: for instance, in [22, §3.2], the assumption that $\lim_{t\to 0}|tX|=1$ is justified by the fact, proved in [29, Thm. 2.3], that this property holds generically—though not always—for finite metric spaces.
@@
Section 2 also contains new examples illustrating just how ‘badly’ continuity can fail. In [29, Thm. 3.8] it was shown that for every real number $\alpha\geq 1$, there exists a finite metric space $X$ such that $\lim_{t\to 0}|tX|=\alpha$, but the question was left open of whether there exists $X$ such that $\lim_{t\to 0}|tX|<1$ or $\lim_{t\to 0}|tX|=+\infty$. That question is not answered here, but in Example 2.6 we show that if one considers arbitrary paths $(X_{s})_{0<s\ll 1}$ in $\operatorname{FMet}^{\circ}$ converging to the one-point space, then $\lim_{s\to 0}|X_{s}|$ can take any value in the extended real line. We also consider certain paths converging to a two-point space $K_{2}$, and describe the associated sequences of magnitude functions. We demonstrate that the pointwise limit of these functions need not display any of the good qualitative features of the magnitude function of $K_{2}$, such as continuity or monotonicity with respect to the scale factor $t$, or meaningful behaviour in the large- or small-scale limit.
@@
An alternative route to a stability theorem for magnitude would be to look for a class of metric spaces such that if $X$ belongs to this class, then magnitude is continuous with respect to the Hausdorff topology on finite subsets of $X$. One good candidate might be the class of positive definite spaces. In [23], Meckes defines a positive definite metric space to be one such that every finite subspace has positive definite similarity matrix, and he proves that magnitude is lower semicontinuous—though not continuous—on the Gromov–Hausdorff space of compact positive definite metric spaces [23, Thm. 2.6]. He also poses the following question:
@@
Since the metric induced by the $\ell^{p}$-norm on $\mathbb{R}^{n}$ is positive definite for $1\leq p\leq 2$ [23, Thm. 3.6], a positive answer to 1.2 would imply in particular that magnitude is continuous for finite subsets of these spaces. This would be especially valuable to know for the purposes of applications, which often concern data embedded into $\mathbb{R}^{n}$ with the Euclidean ($\ell^{2}$) or taxicab ($\ell^{1}$) metric.
Moreover, in the setting of finite-dimensional, positive definite normed spaces such as these, the techniques of Meckes [24] are available, which seem likely to facilitate stronger results on continuity. Indeed, the strongest known result in this direction says that magnitude is continuous on the class of $n$-dimensional compact convex sets in any $n$-dimensional positive definite normed space [20, Thm. 4.15]. (In fact, the statement is a bit stronger than that.) And in Leinster and Meckes [21, Thm. 3.1], the techniques of [24] are used to prove that if $U$ is a finite-dimensional, positive definite normed space then every compact subset $X$ of $U$ satisfies $\lim_{t\to 0}|tX|=1$. (Other relevant remarks by Meckes concerning the behaviour of magnitude in this setting can be found in [24], following the proof of Theorem 1.) We therefore close this introduction by stating a conjecture.
@@
Let $U$ be a finite-dimensional, positive definite normed space. Then the function $X\mapsto|X|$ is continuous with respect to the Hausdorff topology on the set of finite subsets of $U$. In particular, magnitude is stable for finite subsets of Euclidean or taxicab space.
@@ Acknowledgements
This work was partially supported by JSPS KAKENHI JP22K18668, JP23H00081. We are grateful to Sergei O. Ivanov for useful conversations, and to Mark Meckes for valuable comments on the first version of this paper.
@@ 2. Magnitude is nowhere continuous on Gromov–Hausdorff space
    [29, Example 3.4] Let $X=\frac{4}{3}K_{2}+_{1}2K_{3}$. Then $\lim_{t\to 0}|tX|=\nicefrac{{7}}{{6}}$.
@@
Take any $X\in\operatorname{FMet}^{\circ}$. Let $Y$ be a space without the one-point property—for concreteness, we may choose $Y=K_{3}+_{1}2K_{3}$ as in Example 2.4 (1). Pick any points $x\in X$ and $y\in Y$, take $t>0$ and consider the wedge $X\vee tY=X\vee_{x,y}tY$. Observe that $X$ embeds isometrically into $X\vee tY$, and that in the Hausdorff metric on subspaces of the wedge, we have $d(X,X\vee tY)\leq\operatorname{diam}(tY)\to 0$ as $t\to 0$. It follows that $X\vee tY\to X$ in $\operatorname{FMet}$. Meanwhile, for sufficiently small $t>0$ the space $tY$ has magnitude [29, Rmk. 2.1], so, by Lemma 2.3 (1), for sufficiently small $t>0$ the space $X\vee tY$ has magnitude given by $|X\vee tY|=|X|+|tY|-1$. Hence, as $t\to 0$, we have
@@
In [29, Thm. 3.8], it was shown that magnitude’s failure of continuity at the one-point space can be ‘arbitrarily bad’, in the sense that for each real number $\alpha\geq 1$ there exists a finite metric space $X$ with $\lim_{t\to 0}|tX|=\alpha$. So far as we know, no example has yet been found of a finite metric space $X$ such that $\lim_{t\to 0}|tX|<1$ or $\lim_{t\to 0}|tX|=+\infty$. However, if we consider more general paths in $\mathrm{FMet}$ converging to the one-point space, the limiting magnitude can attain any value in the extended reals. A modification of Example 2.4 realizes such paths.
@@
By [29, Thm. 3.8], for every real number $\alpha\geq 1$ there exists a finite metric space $Y_{\alpha}$ such that $\lim_{t\to 0}|tY_{\alpha}|=\alpha$. Fix $\alpha$ and let $W_{t}=K_{2}\vee tY_{\alpha}$; then $\lim_{t\to 0}W_{t}=K_{2}$. For each $s>0$, we have $sW_{t}=sK_{2}\vee stY_{\alpha}$. It follows, by Lemma 2.3 (1), that
@@ 3. Lines in Gromov–Hausdorff space
In [29, §2], the small-scale limit of the magnitude function was studied via this identification of finite metric spaces with equivalence classes of matrices. To adapt the techniques of that paper to the present setting, we will identify lines in Gromov–Hausdorff space with certain equivalence classes of block matrices.
@@
Remark 3.7.
Some readers may prefer to think of a line to $X$ as representing a direction of approach to $X$: lines $(Y_{t})$ and $(Y_{t}^{\prime})$ approach ‘from the same direction’ if $Y_{1}^{\prime}\simeq Y_{t}$ for some $0<t\leq 1$, or vice versa. Our main result (Theorem 4.1) then says that for a generic finite metric space $X$, and a linear sequence approaching $X$ from a generic direction, magnitude behaves continuously. Thus, for continuity of magnitude, the bad directions are rare, and almost all directions are good.
@@ 4. Magnitude behaves continuously for generic limits along lines
hence, $\det(Z_{f,t})$ is a generalized polynomial in $e^{-t}$, in the sense of [29, §2] (but now with real coefficients). It follows that $\det(Z_{f,t})$ vanishes for at most finitely many $t$. In particular, for all sufficiently small $t>0$ we must have $\det(Z_{f,t})\neq 0$, so the expression (4.2) is valid for $0<t\ll 1$. We will compute the denominator and the numerator of (4.2) separately, beginning with the denominator.